OpenStudy (anonymous):

Keith spent half of his allowance going to the movies he washed the family car and earned $7 dollars what is his weekly allowance if he ended with 15 dollars?

OpenStudy (anonymous):

1 week = 7 days, x be his weekly allowance (x+7)-0.5x=15 0.5x = 8 x = 16 So he gets 16 dollars a week. cheers

OpenStudy (anonymous):

but how tho if he ended w/15 dollars?

OpenStudy (anonymous):

easy you work backwards from here, so! if he gets 16 dollars a week, he get extra 7 dollars for washing the car, so total 23, then he went movie which cost him 8dollars because the question states he spend half his allowance which we know 16/2 = 8! so 23-8=$15 left WALA!

OpenStudy (anonymous):

WOW u are smart thank you so much!!!!!!!!!!!!!!!